OpenStudy (anonymous):

cos^2x + 2 cos x + 1 = 0

OpenStudy (anonymous):

Find all solutions

OpenStudy (solomonzelman):

\(\LARGE\color{blue}{ \bf cos^2(x) + 2 cos( x) + 1 = 0 }\) factor just like doing a^2+2a+1=(a=1)^2 \(\LARGE\color{blue}{ \bf ( cos~x+ 1) ^2= 0 }\) take square root of both sides \(\LARGE\color{blue}{ \bf cos~x+ 1= 0 }\) \(\LARGE\color{blue}{ \bf cos~x=-1 }\)